### v1

# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

Abstract
This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for unknown environments. We consider the pixel-wise environment perception problem where one is given 2D range data from LiDAR scans and must label points contextually as map or non-map in the surroundings for visual planning. LIVES classifies incoming 2D scans from the wide Field of View (FoV) LiDAR in unseen environments without prior map information. The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier in real environments. A visual planner takes contextual data from scans and uses this information to plan viewpoints more likely to yield detection of the search target. While conventional frontier based methods for LiDAR and multi sensor exploration effectively map environments, they are not tailored to search for people indoors, which we investigate in this paper. LIVES is baselined against several existing exploration methods in simulation to verify its performance. Finally, it is validated in real-world experiments with a Spot robot in a 20x30m indoor apartment setting. Videos of experimental validation can be found on our project website at https://sites.google.com/view/lives-icra-2024/home.
I INTRODUCTION
Autonomous planning and real-world execution for active search and exploration are receiving significant attention from the robotics community due to their relevance in several scenarios including inspection, surveillance, and Search and Rescue (SAR) [27]. Autonomous ground robots cannot assume reliable map information a priori during active search missions, for example in a disaster response scenario. This demands environment perception tailored for the task of locating any persons in need quickly. Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. However, use of these scans is often limited to the task of localization and mapping. This paper addresses the exploitation of wide Field of View (FoV) LiDAR scans for visual planning. We propose to classify LiDAR scan points as map or non-map in the robot’s surroundings for improved visual search planning.
Several methods exist for pixel-wise classification of 3D LiDAR scans [22, 23, 34]. However, [15] notes that these methods are limited by the high computational requirements to handle to large input size of 3D cloud data. In contrast, real time projection-based methods aim to detect moving obstacles [13] or objects to be avoided in 3D scans, often for autonomous driving applications [18, 5, 20]. Despite parallel improvements in efficiency of planning algorithms for both LiDAR and vision sensors [6, 14, 28, 32] it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search. There exists a gap of real-time pixel-wise classification of LiDAR scans for active visual search tasks. This work addresses this limitation with an online 2D pixel-wise LiDAR scan classifier and context informed visual planner. In particular, classified LiDAR scans are used to generate highly informative plans quickly with a multisensor Frontier Exploration [33] based planner.
Figure 1: Overview of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps (d). Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints. Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2. In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. State of the art visual task planners like [6, 32] account for LiDAR scans without considering contextual information. We provide a visual planner with map and non-map data in the surroundings found in wide FoV LiDAR scans. This is accomplished by learning to distinguish points unlikely to be permanent features in unknown environments. The visual planner gives priority to features of interest under the assumption that the search target belongs to the set of non-map points. This information obviates the need to visually inspect parts of the environment that are likely to be permanent. An overview of the proposed method can be found in Fig. 1.
The contributions of this work can be summarized as follows:
  • A learning strategy for pixel-wise scan classification as map and non-map in unseen, real-world environments
  • A formulation that incorporates classified scan information into a planner for visual search
  • Demonstrate reduced mission completion time in the visual search task over existing planners
  • Deploy and verify the proposed method on a real robotic system in an unseen indoor environment
Sections II, III and IV describe related works, methods an results including baseline comparisons.
II Related Works
II-A Lidar Classification
Indoor pointcloud classification is tackled in 3D [30] using the KPConv architecture [29]. This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [31]. Neither work validates on a real robot platform. Ref. [21] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [30], focused on map building. Finally, [26] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping environments. In contrast LIVES recognizes non-map features.
II-B Next-Best View and Informative Path Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [33, 14]. A common approach uses an active perception formulation to maximize entropy reduction [17, 12]. Ref. [11] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [19] similarly learn an information aware policy offline that proposes waypoints for the robot. A local planner then generates dynamically feasible trajectories based on these suggestions. In contrast to the proposed approach, [11, 19] only provide simulation results. Similar to our method, [36] focuses on the problem of visual search. They propose an end-to-end policy that learns to generate discrete actions that lead a robot to a prescribed viewpoint in the given environment, however, the action space is limited to simple discrete actions and they do not leverage LiDAR information. Ref. [35] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, this work is aimed at small areas and small objects, rather than quickly searching buildings for people.
II-C World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [25] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Similarly, [28] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. A planner uses this to explore rooms fully before continuing global exploration, unlike LIVES that avoids complete exploration to reduce search completion time. Ref. [24] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [10] instead learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations.
III Methods
The overall scan classifier and planning module are described in Alg. 1. The key components are detailed in this section.
Algorithm 1 Map-Free LiDAR Informed Search()
$s_{i},x_{i}$
LiDAR Scan, Robot Pose
$x^{*}_{\text{next}}$ (Next viewpoint)
$s_{i}^{\text{class}}=\text{NeuralNetworkClassifier}(s_{i})$
$\mathcal{M}^{\text{LiDAR}}_{i}\leftarrow\text{LiDARMapUpdate}(s_{i},x_{i},\mathcal{M}^{\text{LiDAR}}_{i-1})$
$\mathcal{M}^{\text{Visual}}_{i}\leftarrow\text{VisualMapUpdate}(s_{i},x_{i},\mathcal{M}^{\text{Visual}}_{i-1})$
$\{\mathbb{C}^{\text{LiDAR}}\}=\text{GetLiDARFrontiers}(\mathcal{M}^{\text{LiDAR}}_{i})$
$\{\mathbb{C}^{\text{Visual}}\}=\text{GetVisualFrontiers}(\mathcal{M}^{\text{Visual}}_{i})$
for $n\leftarrow 1$ to $N$ do $\triangleright$ for each frontier $c_{n}$ in $\mathbb{C}$
     for $j\leftarrow 1$ to $4$ do $\triangleright$ for each viewpoint at $c_{n}$
         $\text{util}_{n,j}=\text{ComputeSampleUtility}(c_{n},x_{i},s_{i}^{\text{class}})$
     end for
end for
$x^{*}_{\text{next}}$leftarrow $\arg\max_{n,j}(\text{util}_{n,j})$
return $x^{*}_{\text{next}}$
III-A Environment
The environment is described as a discrete set of points $\mathcal{E}$. This can be split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, which represent the free and occupied cells in the environment. Both of these sets are initially unknown to the robot, but are assumed to follow a structure inherent in indoor environments (e.g. hallways, planar walls, open rooms and loops).
III-B Ground Truth LiDAR Scan Classification
The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$. In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [9]. LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points. Given map information, scan points can be classified into these three categories as follows.
Let $x_{i}$ denote the pose of the robot, and $s_{i}$ denote observation at time step $t_{i}$. Each observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$. Observations are transformed from robot local frame into the global frame using an affine transformation $T_{i}\in SE(3)$. Finally, let map $\mathcal{M}$ be represented as a set of lines $\{l_{i}\}_{1:n}$.
III-B1 LTF
First, an analytic ray cast is performed [8] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to one of the lines of that static map can be written
$P(p_{i}^{j}|x_{i},M)=\text{exp}\left(-\frac{\text{dist}(T_{i}p_{i}^{j},l_{j})^{2}}{\Sigma_{s}}\right)$ (1)
where $\Sigma_{s}$ is the scalar variance of observations, which comes from sensor accuracy. If Eq. 1 is greater than a threshold, point $p_{i}^{j}$ is classified as a LTF.
III-B2 STF
Remaining points will be classified as STF or DF. Observations at current time $i$, $p_{i}^{j}$, are compared with prior observations at time $k$, $p_{k}^{l}$ to determine correspondence between points in subsequent observations. The likelihood of the remaining points corresponding to the same point as in a previous laserscan is computed as
$P(p_{i}^{j},p^{l}_{k}|x_{i},x_{k})=\text{exp}\left(-\frac{||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||^{2}}{\Sigma_{s}}\right)$ (2)
where $p^{l}_{k}$ is the nearest point from $p_{i}^{j}$ among points which does not belong to LTF at other timesteps, defined as
$p^{l}_{k}=\arg\text{min}||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||$ (3)
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$. In practice, there are a large number of STFs and DFs due to the size of scans, therefore it is important to filter them. First, the pooling operator is used to reduce duplicates within a radius. Second, pooled STFs and DFs within some distance of LTFs are removed to eliminate false positives caused by localization drift. Finally, points inside visually observed regions of the map are removed.
III-C Map-free LiDAR Scan Classification
Figure 2: Pixel-wise LiDAR scan classification model architecture. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN), one for each input. The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$. Similarly to the ground-truth, non-map points are filtered by pooling and then by removing those in previously inspected regions.
III-C1 Dataset and Data Acquisition
The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time. This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places. To ease data acquisition we design a steerable cart platform (see fig. 3) to mimic the robot. Human operators easily maneuver this cart to generate the dataset without strain to the robot. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. Finally, a laptop is used to maintain the pose estimate [9], collecting the data and generating online ground-truth labels for the dataset. This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online.
Figure 3: A figure depicting the cart used for labeled data acquisition and the Spot robot used for deployment of the proposed method. The RealSense provides odometry estimate to the cart platform, required for localization and ground-truth estimation. The spot is equipped with an RGB-D Azure Kinect for detection during the search task.
III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window. The history buffer is then concatenated with the updated pixel-wise exponential weighted average. At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized.
The policy consists of a temporal-convolutional encoder (TCN) for each input. Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each encoder is the same size as the LiDAR scan to be classified. The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function. The resulting LiDAR scan encoding is multiplied element-wise with the label encoding. A threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes. The full model architecture is shown in Fig. 2.
III-D Map Updates
Two global occupancy maps (search maps) are maintained. The first is a LiDAR global search map and the second is a visual global search map. Global search maps are implemented in this work using ROS Costmap2D [2]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$. This process is shown in Fig. 1(d).
III-E Multisensor Frontier Exploration
The planning module seeks to generate a viewpoint at each planning step that results in the highest likelihood to find the target of interest. This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and finally the point with the highest score is fed to the robot for navigation to that point.
III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [33], shown in Fig. 1(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Candidate viewpoints at frontiers are similar to those of [32]. First, a set of vision based viewpoints are generated by clustering frontiers on the visual global search map. Next, a set of LiDAR based viewpoints are generated by clustering frontiers on the LiDAR global search map. Each frontier cluster becomes four candidate viewpoints, facing each direction, to account for the limited FoV of the vision sensor.
III-E2 Viewpoint Selection
Candidate viewpoints are selected based on a computationally efficient heuristic utility function. High value viewpoints are those that are 1) near the robot, 2) allow inspection of many unknown map points and 3) result in inspection of non-map points. This process is shown in Fig. 1(f). The utility functions is a weighted sum of the following:
  • Subtract path distance from robot position to the viewpoint
  • Add a reward for expected number of unknown cells discovered at that viewpoint
  • Add a reward for viewpoints with frontiers near the path to the candidate viewpoint
  • Add a reward for viewpoints that result in inspection of a scaled number non-map points
Scaling parameters on each term are tuned during experiments. Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed on real robot hardware and used in a human environment.
IV-A LiDAR Scan Classification
Figure 4: Ablation studies over policy architecture (top) and injected noise during training (bottom). The inclusion of the label history buffer yields 11.63% higher test accuracy. The policy is robust up to 30% to bit-flipping errors in the label history buffer. The mean accuracy depicts the 5-step moving average.
First we compare the learned classification policy to the ground truth classifier. The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point. If this value is non-negative, the point is classified as a member of $\mathcal{E}^{\text{map}}$. Otherwise the point is classified as a member of $\mathcal{E}^{\text{non-map}}$. We express performance in terms of per-scan accuracy of the classified LiDAR scans from the neural network estimator versus ground truth scan classification from map information. This metric for a classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
$\text{accuracy}=\frac{1}{{n_{i}}}\sum^{n_{i}}_{j=0}\{+1\text{ if }p^{j}_{i}==p^{j,\text{true}}_{i}\text{ else }0\}$ (4)
The final test accuracy of the policy compared to ground truth is $86.19\%\pm 0.03\%$ with no a priori map requirement. We perform two ablation studies, modulating the model architecture and injected noise during training, with results in Fig. 4. The inclusion of the label encoder is responsible for a 11.63% increase in test accuracy. Because the robot moves relatively slowly compared to the policy update frequency, recently predicted labels strongly bias the current estimate toward the correct classification. We fix the history buffer to 9 time steps, corresponding to 1.8 seconds. The history buffer enables the policy to reason about its environment even though no reference map is provided. During training, uniform random noise is injected into the label history buffer to simulate inference-time auto-regressive accumulated errors. The policy is robust up to 30% of injected noise before the policy begins to over-fit and performance degrades. We select a policy trained with 10% injected noise to deploy on the robot. All ablation studies were performed with a fixed seed (0) and initialized with identical weights. We do not tune the random seed.
IV-B Simulation and Baseline Comparison
Simulations are performed using Gazebo-ROS on the Turtlebot [4]. ROS move base is used for navigation [1]. Baselines are: 1) a lightly modified version of [32] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [7], 3) a Rapidly-Exploring Random Tree (RRT) [3], and 4) LIVES given ground-truth scan classifications. The Multisensor baseline is equivalent to LIVES and ground-truth without contextual LiDAR information.
Figure 5: The Apartment simulation environment (top row) in the Easy configuration (a) and Hard Configuration (b). The Office (c) shows the robot starting position as a red arrow. Easy setup involves the search target to the left of the cyan line and Hard involves the search target on the right of the cyan line.
Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
Name Difficulty [7] [3] Exploration [32] Classification (Our Method)
Apartment Easy $112(70\%)$ $72(80\%)$ 36 (100%) 34(100%) 34(100%)
Hard $177(30\%)$ $128(60\%)$ 106 (100%) 92(100%) 90(100%)
Office Easy $70(80\%)$ $54\pm(80\%)$ 34 (100%) 26(100%) 24(100%)
Hard $300(0\%)$ 111 (80%) $141(90\%)$ 99(100%) 105(100%)
TABLE I: Average time (s) for each method to find the object of interest in the given simulation environment.
Two Gazebo world environments are tested with easy and hard setups in each. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig. 5(a-b)). Here the easy setting incorporates fewer objects in the environment and the search target being closer to the robot while hard is defined by more objects in the environment and larger distance to the target. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig. 5(c)). The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method. We see a large gap, however, in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. Overall, the results indicate that the proposed method to inform the frontier planner with non-map information outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
Figure 6: A time lapse of one experiment scenario. In (a) the robot pose is a purple arrow with the visual costmap, global search map, frontier points and non-map points shown. The raw scan data is shown as small white squares. The target object is behind the robot starting position. Figure (b) occurs immediately before a planning step, wherein the marked frontier is selected due to the high occurrence of non-map points nearby. (c) occurs after the planning step and results in detection.
IV-C Hardware Experiments
Experiments are in a 20mx30m indoor apartment setting. The environment is out of distribution as it is not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. The platform used is a Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera. The search target of interest is a small rolling suitcase and YOLOv5 [16] is used for detection of the object. On detection, the exploration node is shut down and the task is complete. We perform several trials over varying initial robot position and target locations. A time lapse of one experiment is shown in Fig. 6 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments can be found on the project website https://sites.google.com/view/lives-icra-2024/home.
V Conclusions
We present a method for visual search in unseen environments using contextual LiDAR information to inform planning. Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%. Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme. Furthermore, the proposed approach is deployed on a real robot in an unseen environment. Experiments certify both the visual search planner and the learning method for map-free scan classification in real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.

### v3

@@
 Abstract
-This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for unknown environments.
-We consider the pixel-wise environment perception problem where one is given 2D range data from LiDAR scans and must label points contextually as map or non-map in the surroundings for visual planning.
-LIVES classifies incoming 2D scans from the wide Field of View (FoV) LiDAR in unseen environments without prior map information.
-The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier in real environments.
-A visual planner takes contextual data from scans and uses this information to plan viewpoints more likely to yield detection of the search target.
-While conventional frontier based methods for LiDAR and multi sensor exploration effectively map environments, they are not tailored to search for people indoors, which we investigate in this paper.
+This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for target search in unknown environments.
+We consider the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings.
+These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks.
+The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier.
+An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target.
+In order to achieve this, we propose a utility function that accounts for traditional metrics like information gain and path cost and also for the additional contextual information from the scan classifier.
 LIVES is baselined against several existing exploration methods in simulation to verify its performance.
-Finally, it is validated in real-world experiments with a Spot robot in a 20x30m indoor apartment setting.
-Videos of experimental validation can be found on our project website at https://sites.google.com/view/lives-icra-2024/home.
+Finally, it is validated in real-world experiments searching for single and multiple targets with a Spot robot in two unseen environments.
+Videos of experimental validation, implementation details and open source code can be found on our project website at https://sites.google.com/view/lives-2024/home.
 I INTRODUCTION
-Autonomous planning and real-world execution for active search and exploration are receiving significant attention from the robotics community due to their relevance in several scenarios including inspection, surveillance, and Search and Rescue (SAR) [27].
-Autonomous ground robots cannot assume reliable map information a priori during active search missions, for example in a disaster response scenario.
-This demands environment perception tailored for the task of locating any persons in need quickly.
+Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [35], target search [1, 12], disaster response [15, 40, 4], agricultural monitoring [13], reconnaissance and surveillance [16, 20, 23].
+Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments.
+This requires planners equipped to generate high quality plans under uncertainty.
+Conventional methods for LiDAR and multi sensor exploration effectively map environments with focus on completeness, rather than being tailored time critical missions, which we investigate in this paper.
 Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings.
-However, use of these scans is often limited to the task of localization and mapping.
-This paper addresses the exploitation of wide Field of View (FoV) LiDAR scans for visual planning.
-We propose to classify LiDAR scan points as map or non-map in the robot’s surroundings for improved visual search planning.
-Several methods exist for pixel-wise classification of 3D LiDAR scans [22, 23, 34].
-However, [15] notes that these methods are limited by the high computational requirements to handle to large input size of 3D cloud data.
+Use of scans is often limited to obstacle avoidance, localization and mapping.
+In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information.
+At the same time, the potential to use additional information from LiDAR scans for online control has grown with access to parallel computing and more sophisticated GPUs.
+While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], ref.
+[17] notes these methods are limited by the high computational requirements to handle to large input size of 3D cloud data.
 In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26].
-Despite parallel improvements in efficiency of planning algorithms for both LiDAR and vision sensors [6, 14, 28, 32] it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search.
-There exists a gap of real-time pixel-wise classification of LiDAR scans for active visual search tasks.
-This work addresses this limitation with an online 2D pixel-wise LiDAR scan classifier and context informed visual planner.
-In particular, classified LiDAR scans are used to generate highly informative plans quickly with a multisensor Frontier Exploration [33] based planner.
-Figure 1: Overview of LIVES.
+There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31].
+Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43].
+There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation.
+This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
+Figure 1: Overview of a planning step of LIVES.
 In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot.
@@
 Finally, the selected viewpoint is sent back to the robot for execution.
-Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments.
-State of the art visual task planners like [6, 32] account for LiDAR scans without considering contextual information.
-We provide a visual planner with map and non-map data in the surroundings found in wide FoV LiDAR scans.
-This is accomplished by learning to distinguish points unlikely to be permanent features in unknown environments.
-The visual planner gives priority to features of interest under the assumption that the search target belongs to the set of non-map points.
+The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target.
+In particular, this method deviates from traditional exploration methods, which focus on completeness.
+Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings.
+We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information.
+We accomplish this using a map based classifier to train a model to recognize features inherent to many structured, indoor environments.
+The perception model feeds the planner those points unlikely to be permanent, non-map points, for use during viewpoint planning.
+These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment.
+Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points.
 This information obviates the need to visually inspect parts of the environment that are likely to be permanent.
@@
 The contributions of this work can be summarized as follows:
-  • A learning strategy for pixel-wise scan classification as map and non-map in unseen, real-world environments
-  • A formulation that incorporates classified scan information into a planner for visual search
+  • A learning strategy for online scan segmentation as map and non-map in unseen, real-world environments
+  • A formulation that incorporates classified scan pixel information into a planner for visual search
   • Demonstrate reduced mission completion time in the visual search task over existing planners
-  • Deploy and verify the proposed method on a real robotic system in an unseen indoor environment
-Sections II, III and IV describe related works, methods an results including baseline comparisons.
+  • Deploy and verify the proposed method on a real robotic system in multiple unseen indoor environments
 II Related Works
@@
 Neither work validates on a real robot platform.
-Ref.
-[21] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [30], focused on map building.
-Finally, [26] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping environments.
-In contrast LIVES recognizes non-map features.
-II-B Next-Best View and Informative Path Planning
+The proposed work instead focus on low computation 2D scan pixel classification.Ref.
+[28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [38], focused on map building rather than contextual labeling for planning.
+Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments.
+In contrast LIVES recognizes non-map features for indoor environments.
+II-B Next-Best View and Multisensor Planning
 Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14].
 A common approach uses an active perception formulation to maximize entropy reduction [19, 10].
+Ref.
+[12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs.
 Ref.
@@
 [22] similarly learn an information aware policy offline that proposes waypoints for the robot.
-A local planner then generates dynamically feasible trajectories based on these suggestions.
-In contrast to the proposed approach, [11, 19] only provide simulation results.
-Similar to our method, [36] focuses on the problem of visual search.
-They propose an end-to-end policy that learns to generate discrete actions that lead a robot to a prescribed viewpoint in the given environment, however, the action space is limited to simple discrete actions and they do not leverage LiDAR information.
+These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information.
 Ref.
 [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas.
-In contrast to LIVES, this work is aimed at small areas and small objects, rather than quickly searching buildings for people.
+In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings for people.
+Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks.
+State of the art multisensor exploration with multiple UAVs [3] leverages the single agent volumetric exploration algorithm presented in [42].
+These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping.
+Ref.
+[27] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments.
+These classifications are utilize by a Pan-Tilt-Zoom camera to perform object detection in Search and Rescue applications.
+These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
 II-C World Prediction Based Planning
@@
 [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored.
+Prediction involves recognizing features common to tunnel networks to improve exploration strategies.
 Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers.
-A planner uses this to explore rooms fully before continuing global exploration, unlike LIVES that avoids complete exploration to reduce search completion time.
+The planner uses this to explore rooms fully before continuing global exploration.
+This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time.
 Ref.
@@
 Prediction guide the robot towards likely exit locations.
+This article describe a novel method for world prediction from 2D scans.
 III Methods
@@
 $p^{l}_{k}=\arg\text{min}||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||$ (3)
-When Eq.
-2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF.
-Remaining points in $p_{i}^{j}$ are classified as DFs.
-The result is classified scan $s^{\text{class}}_{i}$.
-In practice, there are a large number of STFs and DFs due to the size of scans, therefore it is important to filter them.
-First, the pooling operator is used to reduce duplicates within a radius.
-Second, pooled STFs and DFs within some distance of LTFs are removed to eliminate false positives caused by localization drift.
-Finally, points inside visually observed regions of the map are removed.
+When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$.
 III-C Map-free LiDAR Scan Classification
-Figure 2: Pixel-wise LiDAR scan classification model architecture.
-$k$ is the length of the history buffer.
-(a) $[3,k]$ pose history matrix containing $[x,y,\theta]$.
-(b) $[n_{i},k]$ LiDAR range history matrix.
-(c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average.
-(d) The model consists of three temporal-convolutional encoders (TCN), one for each input.
-The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan.
-In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
 The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$.
 Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$.
-Similarly to the ground-truth, non-map points are filtered by pooling and then by removing those in previously inspected regions.
 III-C1 Dataset and Data Acquisition
@@
 This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places.
-To ease data acquisition we design a steerable cart platform (see fig.
-3) to mimic the robot.
+To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot.
 Human operators easily maneuver this cart to generate the dataset without strain to the robot.
@@
 In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online.
-Figure 3: A figure depicting the cart used for labeled data acquisition and the Spot robot used for deployment of the proposed method.
-The RealSense provides odometry estimate to the cart platform, required for localization and ground-truth estimation.
-The spot is equipped with an RGB-D Azure Kinect for detection during the search task.
+Training data can be found on the project website.
 III-C2 Architecture and Training
@@
 At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized.
+In practice there are a large number of non-map points due to the size of scans ($n_{i}=897$), therefore we apply filtering steps like pooling and removing those in previously inspected areas before these points are used for planning.
+Code for training a model can be found on the project website.
+Figure 2: Pixel-wise LiDAR scan classification model architecture.
+$k$ is the length of the history buffer.
+(a) $[3,k]$ pose history matrix containing $[x,y,\theta]$.
+(b) $[n_{i},k]$ LiDAR range history matrix.
+(c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average.
+(d) The model consists of three temporal-convolutional encoders (TCN), one for each input.
+The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan.
+In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
+The full model architecture is shown in Fig.
+2.
 The policy consists of a temporal-convolutional encoder (TCN) for each input.
@@
 The resulting LiDAR scan encoding is multiplied element-wise with the label encoding.
-A threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
+A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
 The policy is trained with mean squared error loss against ground truth classifications.
 Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.
-The full model architecture is shown in Fig.
-2.
 III-D Map Updates
@@
 Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers.
-Candidate viewpoints at frontiers are similar to those of [32].
-First, a set of vision based viewpoints are generated by clustering frontiers on the visual global search map.
-Next, a set of LiDAR based viewpoints are generated by clustering frontiers on the LiDAR global search map.
-Each frontier cluster becomes four candidate viewpoints, facing each direction, to account for the limited FoV of the vision sensor.
+Any group of frontier points beyond a minimum size becomes its own frontier.
+Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering.
+Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.
 III-E2 Viewpoint Selection
@@
 IV Results
-We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed on real robot hardware and used in a human environment.
+We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on real robot hardware in unseen indoor environments.
 IV-A LiDAR Scan Classification
@@
 We perform two ablation studies, modulating the model architecture and injected noise during training, with results in Fig.
-4.
+3.
 The inclusion of the label encoder is responsible for a 11.63% increase in test accuracy.
@@
 Simulations are performed using Gazebo-ROS on the Turtlebot [32].
-ROS move base is used for navigation [1].
+ROS move base is used to navigate to the goal points given from the viewpoint selection methods compared [25].
 Baselines are: 1) a lightly modified version of [42] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [5], 3) a Rapidly-Exploring Random Tree (RRT) [41], and 4) LIVES given ground-truth scan classifications.
-The Multisensor baseline is equivalent to LIVES and ground-truth without contextual LiDAR information.
-Figure 5: The Apartment simulation environment (top row) in the Easy configuration (a) and Hard Configuration (b).
-The Office (c) shows the robot starting position as a red arrow.
-Easy setup involves the search target to the left of the cyan line and Hard involves the search target on the right of the cyan line.
+Ref.
+[42] described a highly efficient planner for multisensor 3D exploration in underwater environments that has also been extended to multi-drone exploration [3].
+The modifications from the original paper are as follows.
+First, the method is projected into 2D.
+Second, candidate viewpoint selection is changed.
+Originally, candidate viewpoints are selected surrounding frontier centroids, but in this work the candidate viewpoints are selected at centroids, with candidate views in four directions.
+This modified version of the multisensor frontier method is equivalent to the proposed approach if our method is not given any contextual information from the scan classifier.
 Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
@@
 Table I: Average time (s) for each method to find the object of interest in the given simulation environment.
+Figure 4: The Apartment simulation environment (top row) in the Easy configuration (a) and Hard Configuration (b).
+The Office (c) shows the robot starting position as a red arrow.
+Easy setup involves the search target to the left of the cyan line and Hard involves the search target on the right of the cyan line.
 Two Gazebo world environments are tested with easy and hard setups in each.
 The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig.
-5(a-b)).
-Here the easy setting incorporates fewer objects in the environment and the search target being closer to the robot while hard is defined by more objects in the environment and larger distance to the target.
+4(a-b)).
+Here the easy setting incorporates fewer objects in the environment while hard is defined by more objects in the environment.
 The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig.
-5(c)).
+4(c)).
 The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position.
@@
 However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method.
-We see a large gap, however, in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant.
+There exists a large gap in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant.
 In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%.
-Overall, the results indicate that the proposed method to inform the frontier planner with non-map information outperforms the baselines across the full spectrum of environments.
+Overall, the results indicate that the proposed method to inform the NBV planner with non-map information outperforms the baselines across the full spectrum of environments.
 We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby.
 These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
-Figure 6: A time lapse of one experiment scenario.
-In (a) the robot pose is a purple arrow with the visual costmap, global search map, frontier points and non-map points shown.
-The raw scan data is shown as small white squares.
+Figure 5: Ground truth maps of the two experimental environments. (a) shows the 15mx38m hallway environment and (b) shows the 20mx30m apartment.
+Figure 6: A time lapse of a simple experiment scenario.
+In (a) the robot pose is a centered on the visual costmap.
+The global search map, frontier points and non-map points are also shown.
 The target object is behind the robot starting position.
-Figure (b) occurs immediately before a planning step, wherein the marked frontier is selected due to the high occurrence of non-map points nearby.
-(c) occurs after the planning step and results in detection.
+The result of the first planning step is also shown.
+The planner takes advantage of the concentration of non-map points behind the robot.
+Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
 IV-C Hardware Experiments
-Experiments are in a 20mx30m indoor apartment setting.
-The environment is out of distribution as it is not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
-The platform used is a Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera.
-The search target of interest is a small rolling suitcase and YOLOv5 [16] is used for detection of the object.
-On detection, the exploration node is shut down and the task is complete.
-We perform several trials over varying initial robot position and target locations.
-A time lapse of one experiment is shown in Fig.
+Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig.
+5.
+The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
+A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized.
+The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object.
+Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics.
+A time lapse of one experiment in the apartment is shown in Fig.
 6 where the search target is situated behind the robot starting position.
@@
 However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby.
+Finally, we compare the performance of the trained classifier in these out of distribution environments to ground truth information.
+The scan classifier accuracy is also evaluated in the two out of distribution experimental environments.
+The results show $84.0168\%$ accuracy in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
+These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments.
 The success of LIVES onboard a real robot verify hypothesis (3).
-Videos of experiments can be found on the project website https://sites.google.com/view/lives-icra-2024/home.
+Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
 V Conclusions
@@
 Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%.
-Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme.
-Furthermore, the proposed approach is deployed on a real robot in an unseen environment.
-Experiments certify both the visual search planner and the learning method for map-free scan classification in real-world scenarios.
+The proposed approach is deployed on a real robot in an unseen environment to certify the learning method and visual planner for real-world scenarios.
 Ongoing work aims to extend the LIVES planner to multiple robots.

### v4

@@
 In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information.
-At the same time, the potential to use additional information from LiDAR scans for online control has grown with access to parallel computing and more sophisticated GPUs.
-While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], ref.
-[17] notes these methods are limited by the high computational requirements to handle to large input size of 3D cloud data.
+At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing.
+While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17].
 In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26].
@@
 The utility functions is a weighted sum of the following:
-  • Subtract path distance from robot position to the viewpoint
-  • Add a reward for expected number of unknown cells discovered at that viewpoint
-  • Add a reward for viewpoints with frontiers near the path to the candidate viewpoint
-  • Add a reward for viewpoints that result in inspection of a scaled number non-map points
+  • Subtract path distance from robot position to viewpoint
+  • Reward expected number of unknown cells discovered
+  • Reward viewpoints with frontiers near the path to the candidate viewpoint
+  • Reward viewpoints that result in inspection of a scaled number non-map points
 Scaling parameters on each term are tuned during experiments.
@@
 Table I: Average time (s) for each method to find the object of interest in the given simulation environment.
-Figure 4: The Apartment simulation environment (top row) in the Easy configuration (a) and Hard Configuration (b).
-The Office (c) shows the robot starting position as a red arrow.
-Easy setup involves the search target to the left of the cyan line and Hard involves the search target on the right of the cyan line.
 Two Gazebo world environments are tested with easy and hard setups in each.
-The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig.
-4(a-b)).
-Here the easy setting incorporates fewer objects in the environment while hard is defined by more objects in the environment.
-The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig.
-4(c)).
+A figure depicting them is shown on the project website.
+The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES.
+The easy setting incorporates fewer objects in the environment while hard has more objects in the environment.
+The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position.
 The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position.
@@
 Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig.
-5.
+4.
 The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
@@
 However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby.
-Finally, we compare the performance of the trained classifier in these out of distribution environments to ground truth information.
-The scan classifier accuracy is also evaluated in the two out of distribution experimental environments.
-The results show $84.0168\%$ accuracy in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
+Finally, we compare the accuracy of the trained classifier in these out of distribution environments to ground truth information.
+The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
 These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments.
@@
 V Conclusions
-We present a method for visual search in unseen environments using contextual LiDAR information to inform planning.
-Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%.
-The proposed approach is deployed on a real robot in an unseen environment to certify the learning method and visual planner for real-world scenarios.
+We present a method for search in unseen environments using contextual information to inform planning.
+Results show LIVES to outperform several baseline methods in varied environments by 10-30%.
+Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme.
+The proposed approach is deployed in an unseen environment to certify the learning method and visual planner for real-world scenarios.
 Ongoing work aims to extend the LIVES planner to multiple robots.

### v5

@@
 While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17].
-In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26].
+Real time projection-based methods instead aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26].
 There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31].
@@
 II Related Works
-II-A Lidar Classification
-Indoor pointcloud classification is tackled in 3D [39] using the KPConv architecture [37].
-This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [38].
-Neither work validates on a real robot platform.
-The proposed work instead focus on low computation 2D scan pixel classification.Ref.
-[28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [39], focused on map building rather than contextual labeling for planning.
-Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments.
-In contrast LIVES recognizes non-map features for indoor environments.
-II-B Next-Best View and Multisensor Planning
+II-A Next-Best View and Multisensor Planning
 Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14].
@@
 [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas.
-In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings for people.
+In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings during a search and rescue scenario.
 Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks.
@@
 These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
-II-C World Prediction Based Planning
+II-B World Prediction Based Planning
 Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first.
@@
 Ref.
-[8] instead learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints.
+[8] learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints.
 Prediction guide the robot towards likely exit locations.
-This article describe a novel method for world prediction from 2D scans.
+II-C Lidar Classification
+Indoor pointcloud classification is tackled in 3D [39] using the KPConv architecture [37].
+This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [38].
+Neither work validates on a real robot platform.
+The proposed work instead focus on low computation 2D scan pixel classification.Ref.
+[28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [39], focused on map building rather than contextual labeling for planning.
+Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments.
+In contrast LIVES recognizes non-map features for indoor environments.
 III Methods
@@
 The utility functions is a weighted sum of the following:
-  • Subtract path distance from robot position to viewpoint
+  • Penalize distance from robot position to viewpoint
   • Reward expected number of unknown cells discovered

### v8

@@
-Learned Contextual LiDAR Informed Visual Search in Unseen Environments
+Fast LiDAR Informed Visual Search in Unseen Indoor Environments
 Abstract
-This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for target search in unknown environments.
-We consider the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings.
+This paper explores the problem of planning for visual search without prior map information.
+We leverage the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings.
 These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks.
-The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier.
+We present LIVES: LiDAR Informed Visual Search, a method aimed at finding objects of interest in unknown indoor environments.
+A robust map-free classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier.
 An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target.
-In order to achieve this, we propose a utility function that accounts for traditional metrics like information gain and path cost and also for the additional contextual information from the scan classifier.
+We propose a utility function that accounts for traditional metrics like information gain and path cost and for the contextual information.
 LIVES is baselined against several existing exploration methods in simulation to verify its performance.
-Finally, it is validated in real-world experiments searching for single and multiple targets with a Spot robot in two unseen environments.
-Videos of experimental validation, implementation details and open source code can be found on our project website at https://sites.google.com/view/lives-2024/home.
+It is validated in real-world experiments with single and multiple search objects with a Spot robot in two unseen environments.
+Videos of experiments, implementation details and open source code can be found at https://sites.google.com/view/lives-2024/home.
 I INTRODUCTION
-Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [35], target search [1, 12], disaster response [15, 40, 4], agricultural monitoring [13], reconnaissance and surveillance [16, 20, 23].
+Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [tan2021comprehensive], target search [balta2017integrated, chen2023hybrid], disaster response [delmerico2017active, tomic2012toward, bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [grocholsky2006cooperative, liang2022reconnaissance, maini2020visibility].
 Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments.
 This requires planners equipped to generate high quality plans under uncertainty.
-Conventional methods for LiDAR and multi sensor exploration effectively map environments with focus on completeness, rather than being tailored time critical missions, which we investigate in this paper.
 Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings.
-Use of scans is often limited to obstacle avoidance, localization and mapping.
-In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information.
+State of the art exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information.
 At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing.
-While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17].
-Real time projection-based methods instead aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26].
-There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31].
-Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43].
-There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation.
-This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
+While several methods exist for 3D LiDAR scan segmentation [qi2017pointnet, qi2017pointnet++, zamorski2020adversarial], these methods are limited by high computational requirements [jhaldiyal2023semantic].
+In contrast, real time projection-based methods often detect moving obstacles to be avoided for autonomous driving applications [chen2022automatic, liu2022less, baur2021slim, mersch2022receding].
+Numerous planners are designed to reduce uncertainty in new environments by exploiting features common to places they are designed for [saroya2020online, ramakrishnan2020occupancy].
+Figure 1: LIVES allows the Spot robot to search for objects quickly in varied indoor environments. Video of experiments can be found at our website.
+Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient, dai2020fast, tao2023seer, vidal2020multisensor], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [wang2019autonomous].
+This work aims to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
 Figure 2: Overview of a planning step of LIVES.
@@
 They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c).
-Separately, LiDAR scans and pose data are used to maintain the global search maps (d).
-Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints.
-Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2.
-In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed.
+Separately, LiDAR scans and pose data are used to maintain the global search maps for the two sensors (d).
+Global search maps are used to generate frontier points, shown in cyan and pink.
+Frontier centroids are shown as white dots and used for sampling.
+Each centroid has four viewpoints scored (e) III-E2.
+In this toy example the left viewpoint in the lower figure in (e) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed.
 Finally, the selected viewpoint is sent back to the robot for execution.
 The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target.
-In particular, this method deviates from traditional exploration methods, which focus on completeness.
+This method deviates from traditional exploration methods, which focus on completeness.
 Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings.
 We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information.
-We accomplish this using a map based classifier to train a model to recognize features inherent to many structured, indoor environments.
-The perception model feeds the planner those points unlikely to be permanent, non-map points, for use during viewpoint planning.
+A map based classifier is used to train a model to recognize features inherent to many indoor environments.
+The perception model feeds the planner the non-permanent points, non-map points, for use during planning.
 These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment.
 Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points.
-This information obviates the need to visually inspect parts of the environment that are likely to be permanent.
+Examples of search environments are shown in Fig.
+1.
 An overview of the proposed method can be found in Fig.
-1.
+2.
 The contributions of this work can be summarized as follows:
@@
   • Demonstrate reduced mission completion time in the visual search task over existing planners
-  • Deploy and verify the proposed method on a real robotic system in multiple unseen indoor environments
 II Related Works
 II-A Next-Best View and Multisensor Planning
-Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14].
-A common approach uses an active perception formulation to maximize entropy reduction [19, 10].
+Exploration is often solved using next-best viewpoint planners like Frontier Exploration [yamauchi1997frontier, dai2020fast].
+A common approach uses an active perception formulation to maximize entropy reduction [julian2014mutual, charrow2015information].
 Ref.
-[12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs.
-Ref.
-[9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas.
-Ref.
-[22] similarly learn an information aware policy offline that proposes waypoints for the robot.
+[chen2023hybrid] combine frontier and sampling based methods while other methods leverage learning to plan [cao2023ariadne, lodel2022look].
 These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information.
 Ref.
-[46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas.
-In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings during a search and rescue scenario.
-Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks.
-State of the art multisensor exploration with multiple UAVs [3] leverages the single agent volumetric exploration algorithm presented in [42].
+[zheng2023system] propose a POMDP for local object search in cluttered 3D environments and extend it with a 2D planner for larger areas, however, their work is aimed at 3D search for small objects, rather than quickly searching large areas.
+Multisensor exploration includes LiDAR information in planning for visual exploration tasks.
+State of the art multisensor exploration with multiple UAVs [best2022resilient] leverages the single agent exploration algorithm presented in [vidal2020multisensor].
 These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping.
 Ref.
-[27] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments.
-These classifications are utilize by a Pan-Tilt-Zoom camera to perform object detection in Search and Rescue applications.
-These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
+[patel2022lidar] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments, which are used by a Pan-Tilt-Zoom camera.
+These works demonstrate efficiency in a number of real world environments, however none of these methods is designed for the environments considered in this article.
 II-B World Prediction Based Planning
-Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first.
+World prediction methods typically learn a model of relevant environments and exploit their structure to explore high-value regions first.
 Ref.
-[33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored.
-Prediction involves recognizing features common to tunnel networks to improve exploration strategies.
-Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers.
-The planner uses this to explore rooms fully before continuing global exploration.
-This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time.
-Ref.
-[31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context.
-Ref.
-[8] learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints.
+[saroya2020online] use topological features from a database of tunnel networks to inform a frontier-based exploration policy.
+Similarly, ref.
+[tao2023seer] use a prediction module to detect semantic objects and classify frontiers, resulting in exploring rooms first before continuing global exploration.
+The proposed method instead aims to ignore low value regions in order to reduce search completion time.
 Prediction guide the robot towards likely exit locations.
 II-C LiDAR Classification
-Indoor pointcloud classification is tackled in 3D [39] using the KPConv architecture [37].
-This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [38].
-Neither work validates on a real robot platform.
+Indoor 3D pointcloud classification is tackled in [thomas2021self] with KPConv [thomas2019kpconv].
+This requires the offline annotation and map building at instantiation and is used for simulated localization and navigation [thomas2022learning].
 The proposed work instead focus on low computation 2D scan pixel classification.Ref.
-[28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [39], focused on map building rather than contextual labeling for planning.
-Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments.
-In contrast LIVES recognizes non-map features for indoor environments.
+[pomerleau2013comparing] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [thomas2021self], focused on map building.
+Finally, [shan2018lego] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar pointsfor mapping outdoor environments.
 III Methods
-The overall scan classifier and planning module are described in Alg. 1. The key components are detailed in this section.
+The overall scan classifier and planning module are described in Fig. 2 and Alg. 1. The key components are detailed in this section.
 Algorithm 1 Map-Free LiDAR Informed Search()
@@
 In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$.
-In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [7].
+In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [biswas2017episodic].
 LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points.
@@
 III-B1 LTF
-First, an analytic ray cast is performed [6] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$.
+First, an analytic ray cast is performed [biswas2012depth] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$.
 Given observations, the probability that points correspond to one of the lines of that static map can be written
@@
 III-C1 Dataset and Data Acquisition
-The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model.
+The dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model.
 The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan.
 $N$ is the total number of data points collected.
-Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time.
-This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places.
+Data acquisition with a robot requires a human expert to control, which is difficult to scale due to mechanical limits of the robot.
 To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot.
-Human operators easily maneuver this cart to generate the dataset without strain to the robot.
+Human operators easily maneuver this cart to obtain the dataset.
 We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification.
-Finally, a laptop is used to maintain the pose estimate [7], collecting the data and generating online ground-truth labels for the dataset.
-This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$.
+A laptop maintains the pose estimate [biswas2017episodic] and collects the data.
 The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz.
@@
 This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels.
-During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window.
-The history buffer is then concatenated with the updated pixel-wise exponential weighted average.
-At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized.
-In practice there are a large number of non-map points due to the size of scans ($n_{i}=897$), therefore we apply filtering steps like pooling and removing those in previously inspected areas before these points are used for planning.
+During inference, the label history buffer is populated by previous estimates from the policy.
+The history buffer is concatenated with the updated pixel-wise exponential weighted average.
+At initialization, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is full.
 Code for training a model can be found on the project website.
 Figure 3: Pixel-wise LiDAR scan classification model architecture.
+This model enables speeding up search by providing contextual information to the planner.
 $k$ is the length of the history buffer.
@@
 The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan.
-In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
+In (e) a threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
 The full model architecture is shown in Fig.
-2.
+3.
 The policy consists of a temporal-convolutional encoder (TCN) for each input.
@@
 Two global occupancy maps (search maps) are maintained.
-The first is a LiDAR global search map and the second is a visual global search map.
-Global search maps are implemented in this work using ROS Costmap2D [24].
+One for the LiDAR and the other for the visual sensor.
+Global search maps are implemented using ROS Costmap2D [costmap].
 All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$.
@@
 This process is shown in Fig.
-1(d).
+2(d).
 III-E Multisensor Frontier Exploration
@@
 III-E1 Candidate Viewpoint Generation
-Frontier points in each of the two global search maps are generated using a process similar to [44], shown in Fig.
-1(e).
+Frontier points in each of the two global search maps are generated using a process similar to [yamauchi1997frontier], shown in Fig.
+2(e).
 Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers.
@@
 Candidate viewpoints are selected based on a computationally efficient heuristic utility function.
-High value viewpoints are those that are 1) near the robot, 2) allow inspection of many unknown map points and 3) result in inspection of non-map points.
+High value viewpoints are those that are near the robot, allow inspection of many unknown map points and result in inspection of non-map points.
 This process is shown in Fig.
-1(f).
-The utility functions is a weighted sum of the following:
-  • Penalize distance from robot position to viewpoint
-  • Reward expected number of unknown cells discovered
-  • Reward viewpoints with frontiers near the path to the candidate viewpoint
-  • Reward viewpoints that result in inspection of a scaled number non-map points
+2(f).
+The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of a scaled number non-map points.
 Scaling parameters on each term are tuned during experiments.
@@
 IV Results
-We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on real robot hardware in unseen indoor environments.
+We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on a real robot in unseen indoor environments.
+Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
+Name Difficulty [bircher2016receding] [rrt] Exploration [vidal2020multisensor] Classification (Our Method)
+Apartment Easy $112(70\%)$ $72(80\%)$ 36 (100%) 34(100%) 34(100%)
+Hard $177(30\%)$ $128(60\%)$ 106 (100%) 92(100%) 90(100%)
+Office Easy $70(80\%)$ $54\pm(80\%)$ 34 (100%) 26(100%) 24(100%)
+Hard $300(0\%)$ 111 (80%) $141(90\%)$ 99(100%) 105(100%)
+Table I: Average time (s) for each method to find the object of interest in the given simulation environments.
 IV-A LiDAR Scan Classification
+We compare the learned classification policy to the ground truth classifier.
+The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point.
+If a value is non-negative, the point belongs to $\mathcal{E}^{\text{map}}$.
+Otherwise the point belongs to $\mathcal{E}^{\text{non-map}}$.
+We express performance in terms of per-scan accuracy of the classified LiDAR scans from the policy vs.
+ground truth from map information.
+This metric for classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
 Figure 4: Ablation studies over policy architecture (top) and injected noise during training (bottom).
@@
 The mean accuracy depicts the 5-step moving average.
-First we compare the learned classification policy to the ground truth classifier.
-The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point.
-If this value is non-negative, the point is classified as a member of $\mathcal{E}^{\text{map}}$.
-Otherwise the point is classified as a member of $\mathcal{E}^{\text{non-map}}$.
-We express performance in terms of per-scan accuracy of the classified LiDAR scans from the neural network estimator versus ground truth scan classification from map information.
-This metric for a classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
 $\text{accuracy}=\frac{1}{{n_{i}}}\sum^{n_{i}}_{j=0}\{+1\text{ if }p^{j}_{i}==p^{j,\text{true}}_{i}\text{ else }0\}$ (4)
@@
 We perform two ablation studies, modulating the model architecture and injected noise during training, with results in Fig.
-3.
+4.
 The inclusion of the label encoder is responsible for a 11.63% increase in test accuracy.
@@
 IV-B Simulation and Baseline Comparison
-Simulations are performed using Gazebo-ROS on the Turtlebot [32].
-ROS move base is used to navigate to the goal points given from the viewpoint selection methods compared [25].
-Baselines are: 1) a lightly modified version of [42] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [5], 3) a Rapidly-Exploring Random Tree (RRT) [41], and 4) LIVES given ground-truth scan classifications.
-Ref.
-[42] described a highly efficient planner for multisensor 3D exploration in underwater environments that has also been extended to multi-drone exploration [3].
-The modifications from the original paper are as follows.
-First, the method is projected into 2D.
-Second, candidate viewpoint selection is changed.
-Originally, candidate viewpoints are selected surrounding frontier centroids, but in this work the candidate viewpoints are selected at centroids, with candidate views in four directions.
-This modified version of the multisensor frontier method is equivalent to the proposed approach if our method is not given any contextual information from the scan classifier.
-Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
-Name Difficulty [5] [41] Exploration [42] Classification (Our Method)
-Apartment Easy $112(70\%)$ $72(80\%)$ 36 (100%) 34(100%) 34(100%)
-Hard $177(30\%)$ $128(60\%)$ 106 (100%) 92(100%) 90(100%)
-Office Easy $70(80\%)$ $54\pm(80\%)$ 34 (100%) 26(100%) 24(100%)
-Hard $300(0\%)$ 111 (80%) $141(90\%)$ 99(100%) 105(100%)
-Table I: Average time (s) for each method to find the object of interest in the given simulation environment.
-Two Gazebo world environments are tested with easy and hard setups in each.
+Simulations are performed in Gazebo with the Turtlebot [tbot].
+Move base is used to navigate to the given viewpoints [navigation-ros].
+Baselines are: 1) a modified Multisensor Frontier Exploration [vidal2020multisensor], 2) the Next-Best View Planner (NBVP) [bircher2016receding], 3) a Rapidly-Exploring Random Tree (RRT) [rrt], and 4) LIVES given ground-truth scan classifications.
+We make the following modifications to ref.
+[vidal2020multisensor] for our environments.
+The method is projected into 2D and viewpoints are sampled at centroids instead of surrounding them.
+This modified multisensor frontier method is equivalent to the proposed approach without segmented scan information.
+Two Gazebo worlds are tested with easy and hard setups in each.
 A figure depicting them is shown on the project website.
 The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES.
-The easy setting incorporates fewer objects in the environment while hard has more objects in the environment.
+The easy setting has fewer objects in the environment while hard has more objects.
 The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position.
-The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position.
+Easy and hard settings are defined by the search object in the same or different half of the environment as the robot starting position.
 Both difficulty levels in each map are repeated 10 times.
@@
 In the easy apartment setting we observe similar performance among the three planners.
-However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method.
-There exists a large gap in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant.
+In the hard setting LIVES and Ground-Truth perform nearly identically and show a 15% improvement with the addition of non-map points compared with the Multisensor method.
+There exists a large gap in the easy office setting with a near 30% improvement using contextual LiDAR information over the Multisensor variant.
 In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%.
-Overall, the results indicate that the proposed method to inform the NBV planner with non-map information outperforms the baselines across the full spectrum of environments.
+The results indicate that LIVES outperforms the baselines across the full spectrum of environments.
 We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby.
 These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
-Figure 4: Ground truth maps of the two experimental environments. (a) shows the 15mx38m hallway environment and (b) shows the 20mx30m apartment.
-Figure 5: A time lapse of a simple experiment scenario.
-In (a) the robot pose is a centered on the visual costmap.
-The global search map, frontier points and non-map points are also shown.
+IV-C Hardware Experiments
+Experiments are in a 20mx30m apartment and a 15mx38m hallway environment.
+These two environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
+A Boston Dynamics Spot is equipped with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera.
+The search object is one or two small rolling suitcases.
+YOLOv5 [yolov5] is used for detection.
+Objects are localized by first converting PointCloud data into the optical frame and then converting to pixel coordinates using camera intrinsics.
+Bounding boxes are then used to obtain the depth information from coinciding pixels.
+A time lapse of a simple trials in the apartment is shown in Fig.
+5.
+The search target is situated behind the robot starting position.
+This example demonstrates the benefit of non-map information.
+In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the candidate on the left hand side.
+Figure 5: A time lapse of a simple scenario.
+The ground truth map (20x30m) is shown in the lower right corner (a).
+This is unavailable to the robot and not included in the training dataset.
+In (a), the global search map, frontier points and non-map points are also shown.
 The target object is behind the robot starting position.
@@
 Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
-IV-C Hardware Experiments
-Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig.
-4.
-The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
-A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized.
-The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object.
-Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics.
-A time lapse of one experiment in the apartment is shown in Fig.
-5 where the search target is situated behind the robot starting position.
-This example demonstrates the benefit of non-map information.
-In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side.
-However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby.
-Finally, we compare the accuracy of the trained classifier in these out of distribution environments to ground truth information.
+Figure 6: A time lapse of the hallway environment (15x38m) experiment scenario.
+In the lower right corner of (a) is the ground truth map with the suitcase locations and non-map points near the robot shown.
+This information is not in the training dataset and not given to the robot.
+In (a) the robot pose is a centered on the visual costmap.
+The first pose selected inspects the non-map points behind the robot.
+In (b) the robot chooses to turn left at the hallway split, recognizing a trash can as a non-map point.
+This correspondence can be seen between the non-map point in (b) appearing in the viewpoint visited by the robot in (c).
+In (d) the robot selects the circled viewpoint, facing to the left.
+This results in viewpoint (e), inspecting the suitcase, which was recognized as a non-map points in (d).
+In (f) the robot has just localized the first suitcase and takes the decision to return to the fork for non-map points found by that candidate viewpoint, leading to the second suitcase.
+(g) shows the overall robot path overlaid on the ground truth map.
+However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby, resulting in detection.
+We compare the accuracy of the trained classifier in these out of distribution environments to ground truth information.
 The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
@@
 Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
+The video demonstrates three trials.
+The first and second take place in the apartment with one and two search objects, respectively.
+In both apartment scenarios, the robot tends to look towards furniture like couches, shelves and tables.
+Similarly, Spot also tends to inspect trash cans and other objects in the hallway.
+This behavior demonstrates the impact of non-map points and suggests the efficacy of the segmentation policy.
+Several key moments in the hallway search scenario are shown in the time lapse in Fig.
+6.
 V Conclusions
-We present a method for search in unseen environments using contextual information to inform planning.
-Results show LIVES to outperform several baseline methods in varied environments by 10-30%.
-Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme.
-The proposed approach is deployed in an unseen environment to certify the learning method and visual planner for real-world scenarios.
-Ongoing work aims to extend the LIVES planner to multiple robots.
+We introduce LIVES, a method to leverage contextual information for search in unseen environments.
+It is shown to outperform baselines by 10-30% in simulation.
+Through ablative studies, we validate our architecture and training approach for map-free scan classification.
+Our experiments confirm the effectiveness of the visual search planner and learning method in real-world scenarios.
+Ongoing work plans to expand to multi-robot applications.
+VI Acknowledgments
+This research was supported in part by NSF Award #2219236 (GCR: Community Embedded Robotics: Understanding Sociotechnical Interactions with Long-term Autonomous Deployments) and Living and Working with Robots, a core research project of Good Systems, a UT Grand Challenge.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation.
+[…]

### v11

@@
 Abstract
-This paper explores the problem of planning for visual search without prior map information.
-We leverage the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings.
-These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks.
-We present LIVES: LiDAR Informed Visual Search, a method aimed at finding objects of interest in unknown indoor environments.
-A robust map-free classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier.
-An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target.
-We propose a utility function that accounts for traditional metrics like information gain and path cost and for the contextual information.
-LIVES is baselined against several existing exploration methods in simulation to verify its performance.
-It is validated in real-world experiments with single and multiple search objects with a Spot robot in two unseen environments.
+This paper details a system for fast visual exploration and search without prior map information.
+We leverage frontier based planning with both LiDAR and visual sensing and augment it with a perception module that contextually labels points in the surroundings from wide Field of View 2D LiDAR scans.
+The goal of the perception module is to recognize surrounding points more likely to be the search target in order to provide an informed prior on which to plan next best viewpoints.
+The robust map-free scan classifier used to label pixels in the robot’s surroundings is trained from expert data collected using a simple cart platform equipped with a map-based classifier.
+We propose a novel utility function that accounts for the contextual data found from the classifier.
+The resulting viewpoints encourage the robot to explore points unlikely to be permanent in the environment, leading the robot to locate objects of interest faster than several existing baseline algorithms.
+Our proposed system is further validated in real-world search experiments for single and multiple search objects with a Spot robot in two unseen environments.
 Videos of experiments, implementation details and open source code can be found at https://sites.google.com/view/lives-2024/home.
 I INTRODUCTION
-Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [tan2021comprehensive], target search [balta2017integrated, chen2023hybrid], disaster response [delmerico2017active, tomic2012toward, bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [grocholsky2006cooperative, liang2022reconnaissance, maini2020visibility].
+Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [nguyen2024grey], search [chen2023hybrid], disaster response [bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [liang2022reconnaissance].
+As mobile robots begin to proliferate in our communities, visual planning capabilities remain fundamental for autonomous agents to achieve success in human environments [zheng2023system].
 Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments.
@@
 Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings.
-State of the art exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information.
-At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing.
-While several methods exist for 3D LiDAR scan segmentation [qi2017pointnet, qi2017pointnet++, zamorski2020adversarial], these methods are limited by high computational requirements [jhaldiyal2023semantic].
-In contrast, real time projection-based methods often detect moving obstacles to be avoided for autonomous driving applications [chen2022automatic, liu2022less, baur2021slim, mersch2022receding].
-Numerous planners are designed to reduce uncertainty in new environments by exploiting features common to places they are designed for [saroya2020online, ramakrishnan2020occupancy].
-Figure 1: LIVES allows the Spot robot to search for objects quickly in varied indoor environments. Video of experiments can be found at our website.
-Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient, dai2020fast, tao2023seer, vidal2020multisensor], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [wang2019autonomous].
-This work aims to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
-Figure 2: Overview of a planning step of LIVES.
-In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot.
-They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c).
-Separately, LiDAR scans and pose data are used to maintain the global search maps for the two sensors (d).
-Global search maps are used to generate frontier points, shown in cyan and pink.
-Frontier centroids are shown as white dots and used for sampling.
-Each centroid has four viewpoints scored (e) III-E2.
-In this toy example the left viewpoint in the lower figure in (e) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed.
-Finally, the selected viewpoint is sent back to the robot for execution.
-The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target.
-This method deviates from traditional exploration methods, which focus on completeness.
-Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings.
+State of the art visual exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information.
+At the same time, the potential to detect additional information in LiDAR scans for online control has grown with advances in parallel computing.
+While methods exist for 3D LiDAR segmentation [qi2017pointnet, zamorski2020adversarial], they are limited by high computational requirements [jhaldiyal2023semantic].
+In contrast, projection-based methods run online but are used to detect moving obstacles for autonomous vehicles [chen2022automatic, mersch2022receding].
+The literature also demonstrates planning under reduced uncertainty in unseen environments by exploiting features common to the places they are designed for [saroya2020online, ramakrishnan2020occupancy].
+Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient], it remains an open problem to enhance performance in autonomous planning and execution [wang2019autonomous].
+To this end, we propose to exploit features common to many indoor environments using a lightweight scan segmentation method for efficient visual search planning.
+The key insight in this work is to exploit contextual information available in wide Field of View LiDAR scans to filter out points unlikely to be the search target.
+This deviates from traditional exploration methods, which focus on completeness.
+Instead, the proposed method prioritizes regions more likely to be the search target in the robot’s surroundings.
 We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information.
 A map based classifier is used to train a model to recognize features inherent to many indoor environments.
-The perception model feeds the planner the non-permanent points, non-map points, for use during planning.
-These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment.
-Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points.
+The perception model feeds the planner the non-permanent points, non-map points.
+These points guide the Next Best View (NBV) planner towards possible search target locations, assumed to be in the set of non-map points in the environment.
 Examples of search environments are shown in Fig.
-1.
-An overview of the proposed method can be found in Fig.
-2.
+LABEL:fig:spot.
 The contributions of this work can be summarized as follows:
-  • A learning strategy for online scan segmentation as map and non-map in unseen, real-world environments
-  • A formulation that incorporates classified scan pixel information into a planner for visual search
-  • Demonstrate reduced mission completion time in the visual search task over existing planners
+  • Achieve significantly faster search indoors than baselines by augmenting multisensor exploration with contextual information from LiDAR scans
+  • A scan segmentation policy that identifies non-permanent points in unseen, real-world environments
+  • A novel frontier-based exploration formulation that accounts for segmented scan information
 II Related Works
-II-A Next-Best View and Multisensor Planning
-Exploration is often solved using next-best viewpoint planners like Frontier Exploration [yamauchi1997frontier, dai2020fast].
-A common approach uses an active perception formulation to maximize entropy reduction [julian2014mutual, charrow2015information].
+The two common approaches to exploration include frontier [yamauchi1997frontier] and information-theoretic sampling methods [charrow2015information].
+Sampling approaches avoid the expensive frontier generation process but require evaluation of every sample, motivating hybrid approaches [dai2020fast].
+Alternatively, [umari2017autonomous] use RRTs to search for frontier points; achieving efficiency using multiple trees and decoupling tree expansion from movement.
+Another approach expands RRTs with accessible space and executes the edge with the highest information gain in a receding horizon fashion [bircher2016receding].
+The key difference between these methods and ours is the addition of segmented scans, which is shown to significantly improve task completion time.
+State of the art multisensor exploration is shown with an underwater vehicle [vidal2020multisensor] and multiple UAVs [best2022resilient].
+They combine frontier generation for two sensors and use a single utility function to score frontier candidates for surface inspection.
+We modify their method to sample candidates at frontier centroids instead of surrounding them.
+This modified method is equivalent to our proposed planner without the use of segmented scan information and represents a third baseline method.
+This baseline clearly demonstrates the positive effect of the proposed contextual information.
+Partially observable markov decision processes (POMDP) are commonly used in visual object search [kaelbling1998planning, zheng2023system].
+The latter presents a platform-generalizable 3D viewpoint planner, capable of reasoning about viewpoints underneath furniture or between appliances.
+Their work includes a 2D planner for multi-room search, however, it is not tailored to search larger spaces for things that may be in plain sight; a skill that offers real-world application in search and rescue.
+Because our goals involve searching larger spaces quickly, the aforementioned exploration algorithms provide more fair baseline comparisons for the proposed system.
+World prediction is used to learn a model of relevant environments and exploit their structure to explore high-value regions first.
 Ref.
-[chen2023hybrid] combine frontier and sampling based methods while other methods leverage learning to plan [cao2023ariadne, lodel2022look].
-These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information.
+[saroya2020online] use topological features like loops or dead-ends from a database of subterranean tunnels to inform a frontier based exploration policy.
+Instead, [ramakrishnan2020occupancy] considers occupancy anticipation from egocentric RGB-D for navigation by an exploration planner.
 Ref.
-[zheng2023system] propose a POMDP for local object search in cluttered 3D environments and extend it with a 2D planner for larger areas, however, their work is aimed at 3D search for small objects, rather than quickly searching large areas.
-Multisensor exploration includes LiDAR information in planning for visual exploration tasks.
-State of the art multisensor exploration with multiple UAVs [best2022resilient] leverages the single agent exploration algorithm presented in [vidal2020multisensor].
-These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping.
-Ref.
-[patel2022lidar] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments, which are used by a Pan-Tilt-Zoom camera.
-These works demonstrate efficiency in a number of real world environments, however none of these methods is designed for the environments considered in this article.
-II-B World Prediction Based Planning
-World prediction methods typically learn a model of relevant environments and exploit their structure to explore high-value regions first.
-Ref.
-[saroya2020online] use topological features from a database of tunnel networks to inform a frontier-based exploration policy.
-Similarly, ref.
-[tao2023seer] use a prediction module to detect semantic objects and classify frontiers, resulting in exploring rooms first before continuing global exploration.
-The proposed method instead aims to ignore low value regions in order to reduce search completion time.
-Prediction guide the robot towards likely exit locations.
-II-C LiDAR Classification
-Indoor 3D pointcloud classification is tackled in [thomas2021self] with KPConv [thomas2019kpconv].
-This requires the offline annotation and map building at instantiation and is used for simulated localization and navigation [thomas2022learning].
-The proposed work instead focus on low computation 2D scan pixel classification.Ref.
-[pomerleau2013comparing] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [thomas2021self], focused on map building.
-Finally, [shan2018lego] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar pointsfor mapping outdoor environments.
+[caley2019deep] learns to predict locations of exits in building with a convolutional neural network using a database of building blueprints.
+Our proposed system leverages a similar principle to exploit the structures inherent to many indoor environments - like planar walls, hallways, rooms and loops - in order to guide the planner.
+We achieve this in a novel way by segmenting scans to identify non-permanent features in the environment.
+Ultimately, the addition of these features are shown to enable our method to find objects of interest more quickly than the examined baselines.
 III Methods
-The overall scan classifier and planning module are described in Fig. 2 and Alg. 1. The key components are detailed in this section.
+The overall scan classifier and planning module are described in Alg. 1 with key components detailed in this section.
 Algorithm 1 Map-Free LiDAR Informed Search()
@@
 III-A Environment
-The environment is described as a discrete set of points $\mathcal{E}$.
-This can be split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, which represent the free and occupied cells in the environment.
-Both of these sets are initially unknown to the robot, but are assumed to follow a structure inherent in indoor environments (e.g.
+The environment is a discrete set of points $\mathcal{E}$, split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, representing free and occupied cells.
+Both sets are initially unknown to the robot, but are assumed to follow a structure inherent to indoor environments (e.g.
 hallways, planar walls, open rooms and loops).
 III-B Ground Truth LiDAR Scan Classification
-The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$.
-In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$.
+Ground truth classification uses map $\mathcal{M}$ to divide $\mathcal{E}^{\text{occ}}$ into $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$.
 In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [biswas2017episodic].
-LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points.
-Given map information, scan points can be classified into these three categories as follows.
-Let $x_{i}$ denote the pose of the robot, and $s_{i}$ denote observation at time step $t_{i}$.
-Each observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$.
-Observations are transformed from robot local frame into the global frame using an affine transformation $T_{i}\in SE(3)$.
-Finally, let map $\mathcal{M}$ be represented as a set of lines $\{l_{i}\}_{1:n}$.
+LTFs represent permanent features while STFs and DFs are non-permanent static and dynamic points, respectively.
+Given $\mathcal{M}$, represented as a set of lines $\{l_{i}\}_{1:n}$, points are classified into these categories as follows.
+Let $x_{i}$ denote robot pose, and $s_{i}$ denote observation at time step $t_{i}$.
+Observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$.
+Observations are transformed from local to global frame using affine transformation $T_{i}\in SE(3)$.
 III-B1 LTF
 First, an analytic ray cast is performed [biswas2012depth] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$.
-Given observations, the probability that points correspond to one of the lines of that static map can be written
+Given observations, the probability that points correspond to the static map can be written:
 $P(p_{i}^{j}|x_{i},M)=\text{exp}\left(-\frac{\text{dist}(T_{i}p_{i}^{j},l_{j})^{2}}{\Sigma_{s}}\right)$ (1)
-where $\Sigma_{s}$ is the scalar variance of observations, which comes from sensor accuracy. If Eq. 1 is greater than a threshold, point $p_{i}^{j}$ is classified as a LTF.
+where $\Sigma_{s}$ is the scalar variance of observations from sensor accuracy. If Eq. 1 is above a threshold, point $p_{i}^{j}$ is an LTF.
 III-B2 STF
-Remaining points will be classified as STF or DF.
+Remaining points are either STFs or DFs.
 Observations at current time $i$, $p_{i}^{j}$, are compared with prior observations at time $k$, $p_{k}^{l}$ to determine correspondence between points in subsequent observations.
-The likelihood of the remaining points corresponding to the same point as in a previous laserscan is computed as
+The likelihood of a point corresponding between time steps is computed as:
 $P(p_{i}^{j},p^{l}_{k}|x_{i},x_{k})=\text{exp}\left(-\frac{||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||^{2}}{\Sigma_{s}}\right)$ (2)
@@
 $p^{l}_{k}=\arg\text{min}||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||$ (3)
-When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$.
+When Eq.
+2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF.
+Remaining points in $p_{i}^{j}$ are classified as DFs.
+The result is classified scan $s^{\text{class}}_{i}$.
+Open source code is published by Biswas et.
+al.
+[biswas2017episodic] at [enml].
 III-C Map-free LiDAR Scan Classification
 The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$.
-Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$.
+Therefore, this work uses supervised learning to train a model to reproduce binary classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$, where $\{p^{j}_{i}\}_{j=1:n_{i},class}\in\{-1,1\}$
 III-C1 Dataset and Data Acquisition
-The dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model.
+We first collect a dataset, however data acquisition with a robot is difficult to scale due to its mechanical limits.
+To ease data acquisition we design an easily maneuverable cart platform (see fig.
+1) to mimic the robot.
+Data is labeled and recorded online as the cart is pushed through eight indoor environments on the university campus that vary in size and shape in attempt to capture generic features found in different indoor environments.
+Training and test maps can be found on the website.
+A laptop with static map information on board the cart receives odometry and scan data from sensors, then uses the method described in the previous section to maintain a pose estimate and collect dataset $\mathcal{D}$, consisting of tuples of data.
 The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan.
+Note that non-map elements in $s_{i}^{\text{class}}$ are value $-1$ and map elements are value $1$.
 $N$ is the total number of data points collected.
-Data acquisition with a robot requires a human expert to control, which is difficult to scale due to mechanical limits of the robot.
-To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot.
-Human operators easily maneuver this cart to obtain the dataset.
-We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification.
-A laptop maintains the pose estimate [biswas2017episodic] and collects the data.
-The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz.
+Figure 1: The cart used for labeled data acquisition and the Spot robot used for deployment.
+The RealSense provides odometry estimate to the cart, required for ground-truth estimation.
+Spot is equipped with an RGB-D Azure Kinect for detection.
+The Velodyne Puck has $n_{i}=897$ and classified scans are recorded at 5Hz.
 A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time.
-In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online.
 Training data can be found on the project website.
 III-C2 Architecture and Training
-We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference.
+In practice, the model only receives raw scan data, pose estimates and previously predicted labels from the robot during operation and must classify pixel-wise the scan online.
+As a result, we formulate pixel-wise classification as a supervised learning problem during training and an auto-regressive problem during inference.
 Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training.
 Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels.
-Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution.
-We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification.
-This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels.
-During inference, the label history buffer is populated by previous estimates from the policy.
-The history buffer is concatenated with the updated pixel-wise exponential weighted average.
+We corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution.
+In order to produce a $k$ length estimated label history input, we take the exponential weighted average of the $k-1$ corrupted pixel-wise classifications to estimate the current timestep’s classification.
+During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k-1$ sliding-window.
+The history buffer is concatenated with the updated pixel-wise exponential weighted average in the same way as was done during training to obtain the $k$ length estimated label history input.
 At initialization, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is full.
 Code for training a model can be found on the project website.
-Figure 3: Pixel-wise LiDAR scan classification model architecture.
-This model enables speeding up search by providing contextual information to the planner.
+Figure 2: Pixel-wise LiDAR scan classification model architecture used to speed up search by providing information to the planner.
 $k$ is the length of the history buffer.
@@
 (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average.
-(d) The model consists of three temporal-convolutional encoders (TCN), one for each input.
+(d) The model consists of three temporal-convolutional encoders (TCN) (i.e.
+pose encoder, scan encoder, and label encoder).
 The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan.
@@
 The full model architecture is shown in Fig.
-3.
-The policy consists of a temporal-convolutional encoder (TCN) for each input.
-Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer.
+2 and consists of a temporal-convolutional encoder (TCN) for each input.
+Each TCN contains a single convolutional layer with scan-wise circular padding and a single linear layer.
 The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$.
 Hyperbolic tangent activations are used for all layers.
-The output of each encoder is the same size as the LiDAR scan to be classified.
-The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function.
-The resulting LiDAR scan encoding is multiplied element-wise with the label encoding.
+The output of each TCN is the same size as the LiDAR scan to be classified ($n_{i}=897$).
+The output of the pose history TCN is summed with the output of the scan encoder as a pose correction before element-wise multiplication with the label encoding, which may flip the sign of the term.
+The result is normalized by applying a hyperbolic tangent function such that each value is $\{-1,1\}$.
 A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
@@
 Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.
+Figure 3: Map updates are performed from observation at each timestep for visual (a) and LiDAR (b) sensors.
+White cells are known free and blue/grey cells are unknown.
+Next, frontiers and their centroids are computed for each of the two sensors (c) and (d).
 III-D Map Updates
 Two global occupancy maps (search maps) are maintained.
-One for the LiDAR and the other for the visual sensor.
-Global search maps are implemented using ROS Costmap2D [costmap].
+One for the LiDAR and the other for the visual sensor (see fig.
+3(a-b)).
+Global search maps are implemented as occupancy grids.
+They are updated at each step from local sensor occupancy grids, which are implemented using the ROS Costmap2D package [costmap].
 All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$.
@@
 Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$.
-This process is shown in Fig.
-2(d).
-III-E Multisensor Frontier Exploration
-The planning module seeks to generate a viewpoint at each planning step that results in the highest likelihood to find the target of interest.
-This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and finally the point with the highest score is fed to the robot for navigation to that point.
+III-E Viewpoint Planning
+The planner sends the viewpoint at each planning step that results in the highest likelihood to find the target of interest.
+This involves picking viewpoints that balance exploration of unknown space in the environment with exploiting non-map points from LiDAR scans.
+This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and best sample is fed to the robot.
 III-E1 Candidate Viewpoint Generation
 Frontier points in each of the two global search maps are generated using a process similar to [yamauchi1997frontier], shown in Fig.
-2(e).
+3(c-d).
 Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers.
@@
 Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering.
-Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.
+Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor (see fig.
+4).
+Figure 4: An overview of the viewpoint sampling process.
+Four viewpoints are considered at each centroid, shown as red triangles.
+The highest scoring viewpoint is shown as a yellow triangle to be sent to the robot.
 III-E2 Viewpoint Selection
-Candidate viewpoints are selected based on a computationally efficient heuristic utility function.
+Candidate viewpoints are scored using the following utility function.
 High value viewpoints are those that are near the robot, allow inspection of many unknown map points and result in inspection of non-map points.
-This process is shown in Fig.
-2(f).
-The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of a scaled number non-map points.
+The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of non-map points.
 Scaling parameters on each term are tuned during experiments.
-Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered.
+Weights are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered.
 The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
 IV Results
-We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on a real robot in unseen indoor environments.
-Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
-Name Difficulty [bircher2016receding] [rrt] Exploration [vidal2020multisensor] Classification (Our Method)
+We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) it reduces detection time compared against existing algorithms, and (3) it can be deployed successfully on a real robot in unseen indoor environments.
+Map Target NBVP RRT MFE Ground Truth Our
+Name Difficulty [bircher2016receding] [umari2017autonomous] [vidal2020multisensor] Classification Method
 Apartment Easy $112(70\%)$ $72(80\%)$ 36 (100%) 34(100%) 34(100%)
@@
 Hard $300(0\%)$ 111 (80%) $141(90\%)$ 99(100%) 105(100%)
-Table I: Average time (s) for each method to find the object of interest in the given simulation environments.
+Table I: Average time in seconds (and success rate) for each method to find the object of interest in simulation environments.
 IV-A LiDAR Scan Classification
@@
 We perform two ablation studies, modulating the model architecture and injected noise during training, with results in Fig.
-4.
+5.
 The inclusion of the label encoder is responsible for a 11.63% increase in test accuracy.
@@
 IV-B Simulation and Baseline Comparison
-Simulations are performed in Gazebo with the Turtlebot [tbot].
-Move base is used to navigate to the given viewpoints [navigation-ros].
-Baselines are: 1) a modified Multisensor Frontier Exploration [vidal2020multisensor], 2) the Next-Best View Planner (NBVP) [bircher2016receding], 3) a Rapidly-Exploring Random Tree (RRT) [rrt], and 4) LIVES given ground-truth scan classifications.
-We make the following modifications to ref.
-[vidal2020multisensor] for our environments.
-The method is projected into 2D and viewpoints are sampled at centroids instead of surrounding them.
-This modified multisensor frontier method is equivalent to the proposed approach without segmented scan information.
-Two Gazebo worlds are tested with easy and hard setups in each.
-A figure depicting them is shown on the project website.
-The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES.
-The easy setting has fewer objects in the environment while hard has more objects.
-The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position.
+Simulations are performed in Gazebo with the Turtlebot and ROS navigation.
+Baselines are: 1) modified Multisensor Frontier Exploration (MFE) as described in section II [vidal2020multisensor], 2) Next-Best View Planner (NBVP) [bircher2016receding], 3) Rapidly-Exploring Random Tree (RRT) [umari2017autonomous], and 4) our method, but informed with ground-truth scan classifications as described in section III-B.
+Figure 6: Environments used for baseline comparison.
+Robot start pose is denoted by a cyan circle.
+(a) The Apartment (20x30m) simulation environment in the hard configuration.
+Easy configuration has fewer objects.
+(b) The Office (25x45m) environment.
+Easy setup involves the search target to the left of the red line and hard involves the search target on the right.
+Two Gazebo worlds are tested with easy and hard setups in each (fig.
+6).
+The apartment (20x30m) examines the impact of non-target objects on the planner, the easy setting with fewer objects.
+The office (25x45m) tests the method in a large environment under significantly varied target position.
 Easy and hard settings are defined by the search object in the same or different half of the environment as the robot starting position.
@@
 Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
-Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods.
-In the easy apartment setting we observe similar performance among the three planners.
-In the hard setting LIVES and Ground-Truth perform nearly identically and show a 15% improvement with the addition of non-map points compared with the Multisensor method.
-There exists a large gap in the easy office setting with a near 30% improvement using contextual LiDAR information over the Multisensor variant.
-In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%.
-The results indicate that LIVES outperforms the baselines across the full spectrum of environments.
-We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby.
+Overall, the results support the proposed approach, reporting 100% success and outperforming baselines across all evaluation settings.
+Further, the nearly identical performance between the proposed approach with both predicted and ground truth labels support the efficacy of the perception module.
+While MFE performed similarly to the proposed approach in the easy apartment, the lack of the labeled scan data slowed task completion by 18-42% across the remaining settings.
+NBVP and RRT baselines struggled in most settings, with the former averaging about 50% success.
+RRT outperformed other baselines in the hard office environment, however, was 20% less successful and moderately slower than the proposed approach.
+The overall results support our method, suggesting the benefits of focusing on non-map points.
+This focus allows the agent to move on when the target is unlikely to be nearby.
 These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
+Figure 7: A toy experimental scenario.
+(a) The first planning step.
+Candidate viewpoints are sampled at yellow disks.
+The robot chooses the next viewpoint at the shown cyan arrow in order to visually inspect the concentration of non-map points behind the robot’s starting pose.
+The robot pose on the ground truth map is shown in the lower right (unknown to the robot, not included in the training dataset).
+(b) The concentration of non-map points is found to be two red chairs and the search object (suitcase).
+This result reinforces that the scan classifier successfully recognizes non-permanent features in unseen indoor environments.
 IV-C Hardware Experiments
 Experiments are in a 20mx30m apartment and a 15mx38m hallway environment.
-These two environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy.
+These two environments are not included in dataset $\mathcal{D}$ used to train the classification policy.
 A Boston Dynamics Spot is equipped with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera.
-The search object is one or two small rolling suitcases.
-YOLOv5 [yolov5] is used for detection.
+We first evaluate the accuracy of the trained classifier in these out of distribution environments to ground truth information.
+This is accomplished by teleoperating the robot while running the ground truth and map-free classifiers in parallel.
+The accuracy in the two out of distribution environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
+These suggests the classifier’s ability to distinguish permanent features common to many indoor environments.
+During experiments, YOLOv5 [yolov5] is used for detection.
 Objects are localized by first converting PointCloud data into the optical frame and then converting to pixel coordinates using camera intrinsics.
 Bounding boxes are then used to obtain the depth information from coinciding pixels.
-A time lapse of a simple trials in the apartment is shown in Fig.
-5.
-The search target is situated behind the robot starting position.
-This example demonstrates the benefit of non-map information.
-In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the candidate on the left hand side.
-Figure 5: A time lapse of a simple scenario.
-The ground truth map (20x30m) is shown in the lower right corner (a).
-This is unavailable to the robot and not included in the training dataset.
-In (a), the global search map, frontier points and non-map points are also shown.
-The target object is behind the robot starting position.
-The result of the first planning step is also shown.
-The planner takes advantage of the concentration of non-map points behind the robot.
-Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
-Figure 6: A time lapse of the hallway environment (15x38m) experiment scenario.
-In the lower right corner of (a) is the ground truth map with the suitcase locations and non-map points near the robot shown.
-This information is not in the training dataset and not given to the robot.
-In (a) the robot pose is a centered on the visual costmap.
-The first pose selected inspects the non-map points behind the robot.
-In (b) the robot chooses to turn left at the hallway split, recognizing a trash can as a non-map point.
-This correspondence can be seen between the non-map point in (b) appearing in the viewpoint visited by the robot in (c).
-In (d) the robot selects the circled viewpoint, facing to the left.
-This results in viewpoint (e), inspecting the suitcase, which was recognized as a non-map points in (d).
-In (f) the robot has just localized the first suitcase and takes the decision to return to the fork for non-map points found by that candidate viewpoint, leading to the second suitcase.
-(g) shows the overall robot path overlaid on the ground truth map.
-However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby, resulting in detection.
-We compare the accuracy of the trained classifier in these out of distribution environments to ground truth information.
-The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment.
-These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments.
-The success of LIVES onboard a real robot verify hypothesis (3).
-Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
-The video demonstrates three trials.
-The first and second take place in the apartment with one and two search objects, respectively.
+A simple experiment is shown in Fig.
+7 and demonstrates the benefits of non-map information in guiding the vision sensor in the robot’s surroundings.
+Figure 8: Paths the robot took to find the objects of interest (suitcases) in the video scenarios 2 and 3.
+(a) Two search objects in the 20x30m apartment environment.
+(b) Two search objects in the 15x38m hallway environment.
+Instead of selecting based on number of unknown cells, contextual information leads the agent to inspect many non-map features nearby, resulting in detection.
+Videos of experimentation, implementation details, code, training data, and all maps can be found on the project website https://sites.google.com/view/lives-2024/home.
+The video shows two experiments in the apartment and one in the hallway.
+Fig.
+8 demonstrates the path of the robot in scenarios 2 and 3.
 In both apartment scenarios, the robot tends to look towards furniture like couches, shelves and tables.
 Similarly, Spot also tends to inspect trash cans and other objects in the hallway.
-This behavior demonstrates the impact of non-map points and suggests the efficacy of the segmentation policy.
-Several key moments in the hallway search scenario are shown in the time lapse in Fig.
-6.
+This behavior demonstrates the impact of non-map points and further supports the efficacy of the segmentation policy.
+In the hallway environment the robot is guided by non-map points at several key forks.
+After first inspecting the concentration of non-map points situated behind the robot, the decisions the robot makes at the two key forks it intersects lead it to quickly find the two suitcases in the environment.
+Overall, the results of hardware experiments on the Spot support hypothesis (3).
 V Conclusions
-We introduce LIVES, a method to leverage contextual information for search in unseen environments.
-It is shown to outperform baselines by 10-30% in simulation.
+We introduce a novel visual planner that leverages contextual information found from wide field-of-view LiDAR scans.
+The presented map-free LiDAR classifier identifies non-permanent points in the environment for the planner’s attention.
+The method is shown to outperform baselines by 10-30% in simulation.
 Through ablative studies, we validate our architecture and training approach for map-free scan classification.
 Our experiments confirm the effectiveness of the visual search planner and learning method in real-world scenarios.
-Ongoing work plans to expand to multi-robot applications.
+Future work plans to leverage the non-map points to generate higher quality maps and expand to multi-robot applications.
 VI Acknowledgments